3825 km one way(point A to point B)
average speed 55 mph
i drive 8 hrs/day
how many days would it takes me to go from point A to point B?
Thank you in advance.

Answer by hamsanash1981@gmail.com(151)

Total Time taken to cover the distance of 3825 at speed 55 kmph = 3825/55 = 765/11
If i drive 8 hrs/ day then the number of days it takes to reach point B
= 765/(11*8)
=8 61/88 days or 8.69 days